I really don't understand this and need help please.

The point (46, 0) lies on the line 2x + 10y = 92. Determine if this point is a solution of the system
2x + 10y = 92
-6x – 2y = -24

Answer by Alan3354(69443)

The point (46, 0) lies on the line 2x + 10y = 92. Determine if this point is a solution of the system
------------------------
Sub 46 for x and 0 for y
2x + 10y = 92
2*46 + 0 = 92 as it says
--------------------
-6x – 2y = -24
-6*46 - 0 <> -24 it's not on that line, not a solution to the system
------------------
If you graph it, you get 2 straight lines. The given point is on one of the line, but not the other.
The solution of the system is the point that's on both lines, where they intersect.
That's the point (1,9)